OpenStudy (anonymous):

What is an equation in slope-intercept form for the line that passes through the points (1, -3) and (3, 1)?

OpenStudy (karatechopper):

slope intercept form y=mx+b

OpenStudy (karatechopper):

slope y1-y2/x1-x2

OpenStudy (whpalmer4):

We need to find the slope and the intercept. Slope is rise/run, or the difference in the y values divided by the difference in the x values. In this case, (-3-1)/(1-3) or (1-(-3))/(3-1), both of which = 2. So now we know that our line is y = mx + b, with m = 2. We plug in either known point to find b.

OpenStudy (whpalmer4):

1=2*3+b, b = -5 checking our work, -3 = 2(1) - 5 and 1 = 2(3)-5

OpenStudy (anonymous):

so the answer would be y=2x-5 right?

OpenStudy (whpalmer4):

Yep, m = 2, b = -5, so y = 2x - 5

OpenStudy (anonymous):

Can you explain better to me how to get the slope?

OpenStudy (anonymous):

cuz when i do it by what my text book says to do i get 4/2

OpenStudy (whpalmer4):

That's right - difference in the y values / difference in the x values (must be in the same order!) so you had points (1,-3) and (3,1). Difference in the y values is -3-1 = -4, or 1-(-3) = 4. Difference in the x values is 1-3 = -2, or 3-1 = 2. Slope is therefore either -4 / -2 = 2, or 4 / 2 = 2. For every 1 point we move on the x axis, we go 2 points on the y axis. We started at (1,-3) and moved to (3,1). That's moving 2 points on the x axis (from 1 to 3), so the y value changes by 2 points for each point on the x axis, or 4. We started at (1,-3), added 2 to the x value and added 4 to the y value, bringing us to (3,1).